2. (10 points) Find a basis for the orthogonal complement of span 0 in RS

linear algebra
Answer #1

Vi = V. 1 A -2 2 2 -2-1 o 4 2 4 2 -5 -1 2 Rg -> R2-QR, 42-5-1 4 -2 -5 -R2 4 . -2 1-16 % - 5/4-1/4 -2 -5 -1 -1-16 -/-5/14 -14RREFAS 10-1 - 6 -5/4 (RREF A) X -| -lla lla -5/4 71/4 500 OC-2-164+ áv=d % -262-544-V4 Oc= Z +4- y +5u @ 5/4/ut Za+ Ztbasis for Orthogonal Complement is 16 16 ka CLfirst we form a matrix A with rows as given vectors.

then we find row reduced echelon form of A.using this we find basis for orthogonal complement of span of given vectors.